Ask your own question, for FREE!
Mathematics 7 Online
OpenStudy (anonymous):

You roll two dice. What is the probability of rolling a 1 or a 6 on the second die, given that you rolled an odd number on the first die? A 6X6 table of dice outcomes will help you to answer this question.

OpenStudy (anonymous):

1/2 * 1/3 = 1/6

OpenStudy (anonymous):

thank you so much :D

OpenStudy (zarkon):

the two die rolls are independent so what you got on the first die has no effect on the second.

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!